Why we have to take x^2  as (1-h)^2 and (1+h)^2.....and how we have to take
 

Grade:11

2 Answers

Arun
4 years ago
Dear student
 
I did not get the question. Please post in elaborated manner.
However,
(1-h)² = 1+h²-2h
(1+h)² = 1+h²+2h
Vikas TU
3 years ago
Dear student 
Question is not clear 
Please attach an image, 
We will happy to  help you 
Good Luck
Cheers